archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Core Status CLOSED Percent Complete Details We are going to perform maintenance this evening on the following DSLAM RAM94 TRU75 BNE75 MOL92 Thee intervention will last 30 minutes at each site Date Friday 10 October 2014 05 09AM Reason for closing Done Comments 3 Comment by OVH Friday 10 October 2014 05 08AM Maintenance on RAM94 has been performed Comment by OVH Friday 10 October 2014 05 09AM Maintenance on

    Original URL path: http://status.ovh.es/?do=details&id=7914 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Backend Core Status CLOSED Percent Complete Details We have a small cut on the optical crown for nra 94 and 93 No impact expected the operator was contacted a team is on hand to locate the break Welding work should start around 17 00 Date Friday 10 October 2014 05 07AM Reason for closing Done Comments 1 Comment by OVH Friday 10 October 2014 05 07AM The fiber loop is

    Original URL path: http://status.ovh.es/?do=details&id=7913 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    ADSL SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details Maintenance work will be carried out tonight on DSLAM of MNE69 and MON69 The intervention will last 30 minutes on each site Date Wednesday 08 October 2014 14 33PM Reason for closing Done Comments 2 Comment by OVH Wednesday 08 October 2014 09 55AM The intervention on MON69 is over we continue with MNE69 Comment by OVH

    Original URL path: http://status.ovh.es/?do=details&id=7896 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details DSLAM and the optical hardware in site SUC94 are no longer responding We will investigate Date Friday 26 September 2014 09 02AM Reason for

    Original URL path: http://status.ovh.es/?do=details&id=7803 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details There is an issue with eligibility webservices and meeting scheduling There is an issue eligibility on ovhtelecom fr as well as ordering We will contact Orange to get a timeline for when the service

    Original URL path: http://status.ovh.es/?do=details&id=7712 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Comment by OVH Friday 29 August 2014 13 10PM Our technician is on site they re investigating Comment by OVH Friday 29 August 2014 13 22PM We can t access the NRA we ll try tomorrow morning Comment by OVH Friday 29 August 2014 13 24PM Our technician is on site Comment by OVH Friday 29 August 2014 13 25PM The electric problem on the optic equipment on the NRA

    Original URL path: http://status.ovh.es/?do=details&id=7602 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Incident Category Backend Core Status CLOSED Percent Complete Details We have a fiber cut between SLO13 and PoP OVH There is no impact on traffic The NRA of SLO13 and SMT13 are no longer secure optical security takes over Date Monday 15 September 2014 04 22AM Reason for closing Done Additional comments about closing Fixed Comments 1 Comment by OVH Monday 15 September 2014 04 22AM The problem is being

    Original URL path: http://status.ovh.es/?do=details&id=7683 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    We will update the modems TG788v2 the night of Wednesday 20th to Thursday 21st of August 2014 A cut off in the connection of a few minutes is expected when you restart the modem on the new firmware This new

    Original URL path: http://status.ovh.es/?do=details&id=7541 (2015-10-15)
    Open archived version from archive



  •